# Real-time Estimation of Bound Water Concentration during Lyophilization with Temperature-based State Observers

### v1

@@ Abstract
Lyophilization (aka freeze drying) has been shown to provide long-term stability for many crucial biotherapeutics, e.g., mRNA vaccines for COVID-19, allowing for higher storage temperature. The final stage of lyophilization, namely secondary drying, entails bound water removal via desorption, in which accurate prediction of bound water concentration is vital to ensuring the quality of the lyophilized product. This article proposes a novel technique for real-time estimation of the residual moisture during secondary drying in lyophilization. A state observer is employed, which combines temperature measurement and mechanistic understanding of heat transfer and desorption kinetics, without requiring any online concentration measurement. Results from both simulations and experimental data show that the observer can accurately estimate the concentration of bound water in real time for all possible concentration levels, operating conditions, and measurement noise. This framework can also be applied for monitoring and control of the residual moisture in other desorption-related processes.
**footnotetext: Corresponding author. Email: braatz@mit.edu
@@ Introduction
Three stages of lyophilization comprise (1) freezing, (2) primary drying, and (3) secondary drying, respectively. In freezing, the product and liquid solvent (usually water) are frozen, in which the free water becomes ice crystals, whereas water bound to the organic material between the crystals (aka bound water) retains its noncrystalline state [4]. Primary drying entails the sublimation of the ice crystals [5]. Subsequently, secondary drying is conducted at higher temperature to remove the bound water via desorption [6, 7]. The stability of a lyophilized product is significantly influenced by the amount of bound water, and so monitoring the concentration of bound water is most important [4]. One of the most common techniques is the Karl Fischer titration, which requires sampling of the vial for offline measurements [8, 9, 4]. To avoid process interruption, some online or non-invasive techniques such as near-infrared (NIR) spectroscopy [10] and tunable diode laser absorption spectroscopy (TDLAS) [11] have been proposed. Detailed discussion of tools for the monitoring of secondary drying can be found in Ref. [4].
@@
Instead of direct measurement, a state observer (aka state estimator, observer, estimator) can be used to estimate states that are not measured [12]; the process is known as state estimation. A well-designed observer can replace expensive and complicated sensors in the system, reducing the total cost and complexity of operation. The principle of state estimation is to combine available measurement data of some states with the physics of a system represented by a mechanistic model, and use that information to estimate the other states that are not measured. A variety of mechanistic models for lyophilization are available [13, 1, 14, 15, 16, 6, 17, 9, 18], which establishes a solid foundation for constructing a reliable observer.
@@
Various state observers have been proposed and successfully implemented in chemical processes [19]. In the context of lyophilization, state observers have been extensively studied and applied to the primary drying step, which aims at estimating the temperature, interface position (amount of ice), and relevant parameters such as the heat transfer coefficient [20, 21, 22, 23, 24, 25, 26, 27]. Besides monitoring, some other applications such as process optimization [28] and control [29, 30] have been demonstrated. Observer design for primary drying is straightforward as it mainly concerns heat transfer associated with sublimation and, in many cases, an observer is not even needed as heat transfer-related quantities, e.g., temperature, can be measured easily. In secondary during, heat transfer and desorption dynamics are coupled, making the observer design more challenging and valuable. Currently, applications of state estimation to secondary drying are very limited. The only literature that proposed a state estimation-like strategy for secondary drying is Ref. [9]; the technique is referred to as a soft sensor which requires measurement of the desorption flux for estimating the residual moisture. The procedure in Ref. [9] does not exploit the mathematical structure of a state observer; the key idea is to iteratively solve the optimization to find the moisture content that matches the measured desorption rate. This technique also requires additional equipment specifically for measuring the desorption flux.
@@
In this article, a new technique is proposed for real-time estimation of bound water concentration during desorption, and is applied to secondary drying in lyophilization. The technique relies on a state observer that estimates the concentration of bound water by using temperature measurement and mechanistic understanding of heat transfer and desorption kinetics. The proposed observer is extensively tested with various simulations and experiments. Since accurate bound water measurement is not trivial and involves complex equipment and procedures [31], our observer is formulated such that the only input required is temperature measurement, which is straightforward and very common in every step of lyophilization [4, 32], allowing for the simplest setup and operation compared to any other methods. The proposed framework can also be easily and systematically extended to other desorption-related processes.
@@
A mechanistic model is an important element in a state observer as it contains the knowledge about the physics of a system. Our model is formulated in the rectangular coordinate system by considering one spatial dimension ($z$) and time ($t$) as pictorially shown in Fig. 1. For lyophilization, 1D modeling is nearly always used because its accuracy is comparable to that of multidimensional modeling while being much computationally cheaper and less complicated, whereas 0D modeling (lumped capacity) is not sufficiently accurate [33].
@@
$\rho C_{p}\dfrac{\partial T}{\partial t}=k\dfrac{\partial^{2}T}{\partial z^{2}}-C_{p,g}\dfrac{\partial(N_{g}T)}{\partial z}+\rho_{d}\Delta H_{s}\dfrac{\partial c_{s}}{\partial t}+Q_{v},\qquad t>0,$ (1)
@@
The additional term $Q_{v}$ in (1) describes the effect of microwave irradiation, which provides volumetric heating to the product [34, 35]. This study focuses on conventional lyophilization, i.e., no microwave, and so this term is set to 0 by default. Nevertheless, it is worth noting that our model and observer are designed to accommodate microwave lyophilization as well (see Supplementary Information). The physics of microwave heating is related to dielectric heating, which is influenced by the electric field strength, the microwave frequency, and the dielectric loss factor of each material; detailed discussion can be found in [17, 36, 35]. Nevertheless, this information is not always readily available for all systems, so we model the microwave heating as a single composite term $Q_{v}$ in this case.
@@
$-k\dfrac{\partial T}{\partial z}(H,t)=h(T(H,t)-T_{b}(t)),\qquad t>0,$ (2)
where $h$ is the heat transfer coefficient at the bottom and $T_{b}(t)$ is the bottom shelf temperature. Typically the shelf temperature initially increases linearly as a function of time,
@@
where $T_{b,0}$ is the initial shelf temperature and $r$ is the temperature ramp-up rate. After reaching the maximum temperature $T_{b,\textrm{max}}$, the shelf temperature is kept constant at that value.***The methods apply for general $T_{b}(t)$. At the top surface, heat transfer is negligible compared to the bottom surface, and thus the boundary condition is
@@
The initial temperature of the frozen region is spatially uniform at $T_{0}$,
$T(x,0)=T_{0},\qquad 0\leq z\leq H.$ (5)
@@
The second part of the model concerns bound water desorption. It has been widely accepted in the literature that the linear driving force model can accurately predict the dynamics of bound water desorption despite being one of the simplest adsorption/desorption models [13, 1, 15, 16, 37, 6, 9, 18]. Hence, the desorption kinetics of bound water is described by
$\dfrac{\partial c_{s}}{\partial t}=-k_{s}c_{s},$ (6)
where $k_{s}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [1, 15, 18]
@@
where $A$ is the frequency factor (aka collision frequency) and $E_{a}$ is the activation energy. This representation is identical to the first-order kinetics for chemical reactions. It has been shown via both experiment and simulation that bound water desorption is negligible during primary drying [15, 16], so the initial concentration of bound water in secondary drying is assumed to be uniform at $c_{s,0}$,
@@
The last part of the model focuses on mass transfer of gas/vapor in the pores of the dried product, which usually consists of water vapor $(w)$ and inert gas (in). The continuity equations, assuming ideal gas behaviors for both components, are [13, 1, 14, 15, 16, 6, 17]
@@
where $p(z,t)$ is the partial pressure, $N(z,t)$ is the mass flux (mass flow rate per cross sectional area), $M$ is the molar mass, $R$ is the gas constant, $\varepsilon$ is the porosity, and the subscripts $w$ and ‘in’ denote the water vapor and inert gas, respectively. The expression for $N$ is usually modeled by the dusty-gas model [13, 14, 15, 16, 6]. Note that the total mass flux $N_{g}$ in (1) is $N_{w}+N_{\textrm{in}}$. The initial conditions for both components are
@@
where $p_{w,0}$ and $p_{\textrm{in}}$ are usually defined by the condenser located downstream of the lyophilizer. The boundary conditions are
@@
$\frac{\partial p_{w}}{\partial x}(H,t)=0,\qquad t>0,$ (15)
$\frac{\partial p_{\textrm{in}}}{\partial x}(H,t)=0,\qquad t>0.$ (16)
The main objective of secondary drying is to remove bound water; hence, the concentration of bound water $c_{s}$ is usually the variable of interest [15, 9, 18, 4]. As can be observed from the model equations, the desorption kinetics are mainly influenced by the temperature $T$, so the concentration of bound water can be predicted accurately with only the energy balance and desorption kinetics equations [18, 38, 33]. As such, without loss of generality, mass transfer equations are omitted from the simulations and observer design. This approach simplifies the model equations, parameter estimation, and observer/control design; this simplification does not significantly affect the accuracy of the process model and state observer (see the results in Section 4.1). Nevertheless, it is important to note that mass transfer equations provide information about pressure, so it could be useful for applications related to pressure control and optimization, which is not the scope of this work.
In this work, the above mechanistic model is simulated numerically. The model equations are spatially discretized using the finite volume method, with the details given in Supplementary Information. The final discretized equations can be written as
@@
where $\mathbf{T}\in\mathbb{R}^{m}$ collects the product temperatures $T_{1},\dots,T_{m}$, $\mathbf{c_{s}}\in\mathbb{R}^{m}$ collects the bound water concentrations $(c_{s,1},\dots,c_{s,m})$, $m$ is the number of grid points in the spatial domain, $\mathbf{F}\in\mathbb{R}^{2m}$ is a nonlinear vector function, and $\mathbf{B}\in\mathbb{R}^{2m\times 2}$ is a matrix. To facilitate the observer design, (17) can be rewritten as
@@
where $\mathbf{F_{T}}\in\mathbb{R}^{m},\mathbf{B_{T}}\in\mathbb{R}^{m\times 2}$ represent the dynamics of the temperature part and $\mathbf{F_{c}}\in\mathbb{R}^{m}$ is the nonlinear function for the concentration part. The finite volume method transforms the original partial differential equations (PDEs) into a system of ordinary differential equations (ODEs). The final ODEs (20) and (21) can be integrated by commercial ODE solvers, in which ode15s in MATLAB is used in this work. This technique is known as the method of lines [39].
Lastly, define the average temperature and average concentration,
@@
A state observer (aka state estimator, observer, estimator) is a tool in control theory used for reconstructing the unmeasured states given the available measurements and mechanistic understanding of a system; the process is referred to as state estimation. Those unmeasured states could be internal states that cannot be measured or states that are difficult to measure. Measuring the temperature during lyophilization is relatively simple and accurate, so we design a state observer that uses the temperature measurement to estimate the concentration of bound water, the most important process variable in secondary drying.
@@
$\dfrac{d\hat{\mathbf{T}}}{dt}=\mathbf{F_{T}(\hat{T},\hat{c}_{s})}+\mathbf{B_{T}u}+\mathbf{L_{T}}(\hat{\mathbf{T}}-\mathbf{T}),$ (29)
$\dfrac{d\mathbf{\hat{c}_{s}}}{dt}=\mathbf{F_{c}(\hat{T},\hat{c}_{s})}+\mathbf{B_{c}u}+\mathbf{L_{c}}(\hat{\mathbf{T}}-\mathbf{T}).$ (30)
@@
where $L_{T},L_{c}$ are the real scalars and $\mathbf{J}_{m}$ is an $m$$\times$$m$ matrix of ones. This parameterization suggests that the temperature measurement at each location contributes equally to the observer, leaving only two degrees of freedom for the design, which are the values of $L_{T}$ and $L_{c}$. A well-designed observer should converge the estimated states converge to the true states fast compared to the time scale of the process. The convergence can be evaluated via the estimation errors defined as
$\mathbf{e_{T}}=|\hat{\mathbf{T}}-\mathbf{T}|,$ (33)
$\mathbf{e_{c}}=|\mathbf{\hat{c}_{s}}-\mathbf{c_{s}}|,$ (34)
@@
The state observer proposed in Section 3.1 receives spatially distributed temperature measurement ${\mathbf{T}}$ and provides estimates of both temperature $\mathbf{\hat{T}}$ and concentration $\mathbf{\hat{c}_{s}}$ in real time. Although having a spatially distributed temperature measurement is feasible using current thermal imaging sensors, traditional lyophilization systems do not have such sensors. For example, a thermocouple used for temperature measurement is usually in contact with the bottom of the product, and so only the bottom temperature is available [6, 4]. Therefore, we propose an alternative state observer for this scenario.
@@
$\dfrac{d\hat{\mathbf{T}}}{dt}=\mathbf{F_{T}(\hat{T},\hat{c}_{s})}+\mathbf{B_{T}u}+\mathbf{L_{T}}(\hat{T}_{p}-T_{p}),$ (37)
$\dfrac{d\mathbf{\hat{c}_{s}}}{dt}=\mathbf{F_{c}(\hat{T},\hat{c}_{s})}+\mathbf{B_{c}u}+\mathbf{L_{c}}(\hat{T}_{p}-T_{p}),$ (38)
where $\mathbf{L_{T}}\in\mathbb{R}^{m}$ and $\mathbf{L_{c}}\in\mathbb{R}^{m}$ are the observer gains. Similarly, we can parameterize $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$ using the vector of ones (instead of the matrix of ones as used for the original state observer) such that the only design parameters are the real scalars $L_{T}$ and $L_{c}$. It is worth noting that, since the observer structure and equations are different, the values of $L_{T}$ and $L_{c}$ used for the state observer and modified observers are also different, as shown in the Implementation to Real Systems section.
Most of the results in this article are based on the original state observer as it uses the most complete measurement information, with some results and discussion on the modified observer in the Implementation to Real Systems section.
@@
The structure of the state observer for estimating the concentration of bound water is illustrated in Fig. 2. The proposed observer has the physics of heat transfer and bound water desorption embedded in $\mathbf{F_{T}}$ and $\mathbf{F_{c}}$, which is given by the mechanistic model. The temperature measurement is fed to the observer terms represented by the observer gains $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$. The observer combines the information from the mechanistic model and temperature measurement to converge the estimated bound water concentration to the true values without the need for concentration measurement.
@@
Figure 2: (A) Schematic diagram of the typical lyophilization process. In secondary drying, a glass vial which contains the dried product is heated such that the bound water is desorbed and removed from the product at the top. The mechanistic model is used to describe heat transfer and desorption dynamics. (B) Structure of the proposed state observer. The observer receives the measured temperature and the control input (i.e., the shelf temperature and microwave power) at time step $t$, and estimates the temperature and concentration at the next time step. To initialize the observer at $t=0$, the estimated temperature can be set to the measured value. The initial concentration is completely unknown due to no measurement, so it can be set to some realistic values obtained from the literature or previous experiment.
@@
The most important consideration for observer design is to ensure that the estimated states converge to the true states fast compared to the time scale of the process; i.e., the estimation error converges to zero. In addition, the observer should be designed to be insensitive to measurement noise, that is, that the estimated states are not significantly polluted by measurement noise. The observer gains $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$ play an important role in the overall performance of the observer. Thus, observer design entails selection of the values of these gains. Instead of attempting to search over all elements of the full observer gain matrices $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$, we show in Sections 3.1 and 3.2 that $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$ be parameterized and rewritten as a product of the real scalar $L_{T}$ or $L_{c}$ and the matrix of ones, so the only design parameters are the real scalars $L_{T}$ and $L_{c}$.
[…]
This two-step procedure is simple and practical as it requires only simulations and experimental data, without any detailed mathematical analysis. This technique is generally sufficient for observer design in many applications. In addition to this technique, a more systematic way of simulation-based observer design is to analyze the mathematical structure of the observer, in which more useful insights into the observer dynamics can be extracted. To begin this analysis, we first define the Jacobian of the nonlinear function $\mathbf{F}$ in (17) as
@@
where $\mathbf{F^{\prime}}$ can be calculated analytically or numerically. One of the simplest but efficient ways of analyzing a nonlinear state observer is to approximate the nonlinear parts with the linear equations, i.e., linearization. Linearization of the model equation (17) results in

### v2

@@ Abstract
Lyophilization (aka freeze drying) has been shown to provide long-term stability for many crucial biotherapeutics, e.g., mRNA vaccines for COVID-19, allowing for higher storage temperature. The final stage of lyophilization, namely secondary drying, entails bound water removal via desorption, in which accurate prediction of bound water concentration is vital to ensuring the quality of the lyophilized product. This article proposes a novel technique for real-time estimation of the bound water concentration during secondary drying in lyophilization. A state observer is employed, which combines temperature measurement and mechanistic understanding of heat transfer and desorption kinetics, without requiring any online concentration measurement. Results from both simulations and experimental data show that the observer can accurately estimate the concentration of bound water in real time for all possible concentration levels, operating conditions, and measurement noise. This framework can also be applied for monitoring and control of the residual moisture in other desorption-related processes.
**footnotetext: Corresponding author. Email: braatz@mit.edu****footnotetext: © 2024. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/.
@@ Introduction
Three stages of lyophilization comprise (1) freezing, (2) primary drying, and (3) secondary drying, respectively. In freezing, the product and liquid solvent (usually water) are frozen, in which the free water becomes ice crystals, whereas the bound water retains its liquid state and is adsorbed to the organic material between the ice crystals [4]. In primary drying, the free water (in the form of ice crystals) is removed via sublimation [5]. Subsequently, secondary drying is conducted at higher temperature to remove the bound water via desorption [6]. The stability of a lyophilized product is significantly influenced by the amount of bound water, and so monitoring the concentration of bound water is highly important [4]. One of the most common techniques is the Karl Fischer titration, which requires sampling of the vial for offline measurements [7, 4]. To avoid process interruption, some online or non-invasive techniques such as near-infrared (NIR) spectroscopy [8] and tunable diode laser absorption spectroscopy (TDLAS) [9] have been proposed. Detailed discussion of tools for the monitoring of secondary drying can be found in Ref. [4].
@@
Instead of direct measurement, a state observer (aka state estimator, observer, estimator) can be used to estimate states that are not measured [10]; the process is known as state estimation. A well-designed observer can replace expensive and complicated sensors in the system, reducing the total cost and complexity of operation. The principle of state estimation is to combine available measurement data of some states with the physics of a system represented by a mechanistic model, and use that information to estimate the unmeasured states. A variety of mechanistic models for lyophilization are available [11, 1, 12, 6, 13, 14, 15, 16, 17], which establishes a solid foundation for constructing a reliable observer.
@@
Various state observers have been proposed and successfully implemented in chemical processes [18]. In the context of lyophilization, state observers have been extensively studied and applied to the primary drying step, which aims at estimating the temperature, interface position (amount of ice), and relevant parameters such as the heat transfer coefficient [19, 20, 21, 22, 23, 24, 25, 26]. Besides monitoring, some other applications such as process optimization [27] and control [28, 29] have been demonstrated. However, applications of state estimation to secondary drying are very limited. The only literature that proposed a state estimation-like strategy for secondary drying is Ref. [16]; the technique is referred to as a soft sensor that requires measurement of the desorption flux for estimating the residual moisture. The procedure described in the aforementioned work does not exploit the mathematical structure of a state observer; the key idea is to iteratively solve the optimization to find the moisture content that matches the measured desorption rate. This technique requires a set of equipment for measuring the desorption flux, which is generally available in pilot-scale lyophilizers.
@@
In this article, a new technique is proposed for real-time estimation of bound water concentration during desorption, and is applied to the secondary drying step in lyophilization. The technique relies on a state observer that estimates the concentration of bound water by using the temperature measurement and mechanistic understanding of heat transfer and desorption kinetics. The proposed observer is extensively tested with various simulations and experiments. Since accurate bound water measurement is not trivial and usually involves complex equipment and procedures [30], our observer is formulated such that the only input required is temperature measurement, which is straightforward and very common in every step of lyophilization [4, 31], allowing for the simplest setup and operation compared to any other methods. The proposed framework can also be easily and systematically extended to other desorption-related processes.
This article is organized as follows. Section 2 describes the lyophilization system and derives the corresponding mechanistic model. Section 3 discusses the design and implementation of our state observers. Section 4 presents several simulation and experimental results to demonstrate the performance of the observers. Finally, Section 5 summarizes the study.
@@
A mechanistic model is an important element in a state observer as it contains the knowledge about the physics of a system. This work considers lyophilization of unit does, in which the product is introduced into a number of vials prior to being lyophilized. Our model is formulated in the rectangular coordinate system by considering one spatial dimension ($z$) and time ($t$) as pictorially shown in Fig. 1. For the drying steps in lyophilization, 1D modeling is nearly always used because its accuracy is comparable to that of multidimensional modeling while being much computationally cheaper and less complicated, whereas 0D modeling (lumped capacity) is sometimes not sufficiently accurate [32].
@@
$\rho C_{p}\dfrac{\partial T}{\partial t}=k\dfrac{\partial^{2}T}{\partial z^{2}}-C_{p,g}\dfrac{\partial\left(N_{g}T\right)}{\partial z}+\rho_{d}\Delta H_{s}\dfrac{\partial c_{s}}{\partial t}+Q_{v},\qquad t>0,$ (1)
@@
The additional term $Q_{v}$ in (1) describes the effect of microwave irradiation, which provides volumetric heating to the product [33, 34]. This study focuses on conventional lyophilization, i.e., no microwave, and so this term is set to 0 by default. Nevertheless, it is worth noting that our model and observer are designed to accommodate microwave lyophilization as well (see Section 4.6). The physics of microwave heating is related to dielectric heating, which is influenced by the electric field strength, the microwave frequency, and the dielectric loss factor of each material; detailed discussion can be found in Refs. [15, 35, 34]. Nevertheless, this information is not always readily available for all systems, so we model the microwave heating as a single composite term $Q_{v}$ in this case.
@@
$-k\dfrac{\partial T}{\partial z}(H,t)=h\left(T(H,t)-T_{b}(t)\right),\qquad t>0,$ (2)
where $h$ is the heat transfer coefficient at the bottom and $T_{b}(t)$ is the bottom shelf temperature. Typically, the shelf temperature increases linearly as a function of time,
@@
where $T_{b,0}$ is the initial shelf temperature and $r$ is the temperature ramp-up rate. After reaching the maximum temperature $T_{b,\textrm{max}}$, the shelf temperature is kept constant at that value.******The methods apply for general $T_{b}(t)$. At the top surface, heat transfer is negligible compared to the bottom surface, and thus the boundary condition is
@@
The initial temperature of the dried region is assumed to be spatially uniform at $T_{0}$,
$T(z,0)=T_{0},\qquad 0\leq z\leq H.$ (5)
@@
The second part of the model concerns the desorption of bound water. It has been widely accepted in the literature that the linear driving force model can accurately predict the dynamics of bound water desorption despite being one of the simplest adsorption/desorption models [11, 1, 6, 13, 36, 14, 16, 17]. Hence, the desorption kinetics of bound water is described by
$\dfrac{\partial c_{\textrm{s}}}{\partial t}=k_{s}(c^{*}_{\textrm{s}}-c_{\textrm{s}}),$ (6)
where $c^{*}_{\textrm{s}}$ is the equilibrium concentration of bound water and $k_{s}$ is the rate constant for desorption that exhibits Arrhenius temperature dependence [1, 6, 17]
@@
where $A$ is the frequency factor (aka collision frequency), $E_{a}$ is the activation energy, and $R$ is the gas constant. It is very common in the literature to set $c^{*}_{\textrm{s}}=0$. This simplification produces practically small errors as shown in Ref. [6] and avoids the need for equilibrium data and detailed knowledge about the solid structure [17]. Thus, the final equation is
$\dfrac{\partial c_{s}}{\partial t}=-k_{s}c_{s}.$ (8)
The initial concentration of bound water in secondary drying is assumed to be uniform at $c_{s,0}$,
@@
Mathematically, the main difference between the original equation (6) and simplified version (8) is that the final concentration approaches the equilibrium value for the former and approaches 0 for the latter. This work relies on the simplified equation (8) due to its practicality and adequate accuracy, but the proposed state observer and relevant framework can be applied to both cases because the overall mathematical structure does not change. The original equation (6) and equilibrium data might be needed when there is evidence showing that the equilibrium concentration is significantly high or strict control at the very low concentration region is required.
The last part of the model focuses on the mass transfer of gas/vapor in the pores of the dried product, which usually consists of water vapor $(w)$ and inert gas (in). The continuity equations, assuming ideal gas behaviors for both components, are [11, 1, 12, 6, 13, 14, 15]
@@
where $p(z,t)$ is the partial pressure, $N(z,t)$ is the mass flux (mass flow rate per cross sectional area), $M$ is the molar mass, $\varepsilon$ is the porosity, and the subscripts $w$ and ‘in’ denote the water vapor and inert gas, respectively. The expression for $N$ is usually modeled by the dusty-gas model [11, 12, 6, 13, 14]. Note that the total mass flux $N_{g}$ in (1) is $N_{w}+N_{\textrm{in}}$. The initial conditions for both components are
@@
where $p_{w,0}$ and $p_{\textrm{in,0}}$ are usually defined by the condenser located downstream of the lyophilizer. The boundary conditions are
@@
$\frac{\partial p_{w}}{\partial z}(H,t)=0,\qquad t>0,$ (16)
$\frac{\partial p_{\textrm{in}}}{\partial z}(H,t)=0,\qquad t>0.$ (17)
The main objective of secondary drying is to remove bound water; hence, the concentration of bound water $c_{s}$ is usually the variable of interest [6, 16, 17, 4]. The concentration of bound water can be predicted accurately with the energy balance and desorption kinetics equations as described and shown in Refs. [17, 37, 32], and so recent models usually omit the mass transfer of gas/vapor in the pores of the dried product, i.e., the last part of the model (10)–(17). This approach simplifies the model equations, parameter estimation, and observer/control design, without significant loss in accuracy.
In this work, the above mechanistic model is simulated numerically. The model equations are spatially discretized using the finite volume method as explained in Appendix A. The final discretized equations can be written as
@@
where $\mathbf{T}\in\mathbb{R}^{m}$ collects the product temperatures ($T_{1},\dots,T_{m}$), $\mathbf{c_{s}}\in\mathbb{R}^{m}$ collects the bound water concentrations $(c_{s,1},\dots,c_{s,m})$, $m$ is the number of grid points in the spatial domain, $\mathbf{F}\in\mathbb{R}^{2m}$ is the nonlinear vector function, and $\mathbf{B}\in\mathbb{R}^{2m\times 2}$ is the corresponding matrix. To facilitate the observer design, (18) can be rewritten as
@@
where $\mathbf{F_{T}}\in\mathbb{R}^{m},\mathbf{B_{T}}\in\mathbb{R}^{m\times 2}$ represent the dynamics of the temperature part and $\mathbf{F_{c}}\in\mathbb{R}^{m},\mathbf{B_{c}}\in\mathbb{R}^{m\times 2}$ represent the dynamics of the concentration part. The finite volume method transforms the original partial differential equations (PDEs) into a system of ordinary differential equations (ODEs). The final ODEs (21) and (22) can be integrated by typical ODE solvers, in which MATLAB’s ode15s is used in this work. This technique is known as the method of lines [38]. Lastly, define the average temperature and average concentration,
@@
A state observer (aka state estimator, observer, estimator) is a tool in control theory that is used for reconstructing the unmeasured states given the available measurements and mechanistic understanding of a system; the process is referred to as state estimation. Those unmeasured states could be internal states that cannot be measured or states that are difficult to measure. Measuring the temperature during lyophilization is relatively simple and accurate, so we design a state observer that uses the temperature measurement to estimate the concentration of bound water, the most important process variable in secondary drying.
@@
$\dfrac{d\hat{\mathbf{T}}}{dt}=\mathbf{F_{T}\left(\hat{T},\hat{c}_{s}\right)}+\mathbf{B_{T}u}+\mathbf{L_{T}}\left(\hat{\mathbf{T}}-\mathbf{T}\right),$ (30)
$\dfrac{d\mathbf{\hat{c}_{s}}}{dt}=\mathbf{F_{c}\left(\hat{T},\hat{c}_{s}\right)}+\mathbf{B_{c}u}+\mathbf{L_{c}}\left(\hat{\mathbf{T}}-\mathbf{T}\right).$ (31)
@@
where $L_{T},L_{c}$ are the real scalars and $\mathbf{J}_{m}$ is an $m$$\times$$m$ matrix of ones. This parameterization suggests that the temperature measurement at each location contributes equally to the observer, leaving only two degrees of freedom for the design: $L_{T}$ and $L_{c}$. A well-designed observer should converge the estimated states to the true states fast compared to the time scale of the process. The convergence can be evaluated via the estimation errors defined as
$\mathbf{e_{T}}=\left|\hat{\mathbf{T}}-\mathbf{T}\right|,$ (34)
$\mathbf{e_{c}}=\left|\mathbf{\hat{c}_{s}}-\mathbf{c_{s}}\right|,$ (35)
@@
The state observer proposed in Section 3.1 receives spatially distributed temperature measurement ${\mathbf{T}}$ and provides estimates of both temperature $\mathbf{\hat{T}}$ and concentration $\mathbf{\hat{c}_{s}}$ in real time. Although spatially distributed temperature measurement can be obtained using thermal imaging sensors, traditional lyophilization systems do not have such sensors. For example, a thermocouple used for temperature measurement is usually in contact with the bottom of the product, and so only the bottom temperature is available [14, 4]. Therefore, we propose an alternative state observer for this scenario.
@@
$\dfrac{d\hat{\mathbf{T}}}{dt}=\mathbf{F_{T}\left(\hat{T},\hat{c}_{s}\right)}+\mathbf{B_{T}u}+\mathbf{L_{T}}\left(\hat{T}_{p}-T_{p}\right),$ (38)
$\dfrac{d\mathbf{\hat{c}_{s}}}{dt}=\mathbf{F_{c}\left(\hat{T},\hat{c}_{s}\right)}+\mathbf{B_{c}u}+\mathbf{L_{c}}\left(\hat{T}_{p}-T_{p}\right),$ (39)
where $\mathbf{L_{T}}\in\mathbb{R}^{m}$ and $\mathbf{L_{c}}\in\mathbb{R}^{m}$ are the observer gains. Similarly, we can parameterize $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$ can be parameterized using the vector of ones (instead of the matrix of ones as used for the original state observer) such that the only design parameters are the real scalars $L_{T}$ and $L_{c}$.******Since the observer structure and equations are different, the values of $L_{T}$ and $L_{c}$ used for the state observer and modified observers are also different.
Instead of bottom temperature measurement, some systems may have sensors installed at the top, and hence the temperature at the top surface is measured [40, 39]. In such cases, a similar procedure introduced in this section can be applied.
Most of the results in this article are based on the original state observer as it uses the most complete measurement information, with some results and discussion on the modified observer in Section 4.5.
@@
The structure of the state observer for estimating the concentration of bound water is illustrated in Fig. 2. The proposed observer has the physics of heat transfer and bound water desorption embedded in $\mathbf{F_{T}}$ and $\mathbf{F_{c}}$, which is given by the mechanistic model. The temperature measurement is fed to the observer terms represented by the observer gains $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$. The observer combines the information from the mechanistic model and temperature measurement to converge the estimated concentration to the true values without the need for concentration measurement.
@@
Figure 2: (A) Schematic diagram of the typical lyophilization process. In secondary drying, a glass vial that contains the dried product is heated such that the bound water is desorbed and removed from the product at the top. The mechanistic model is used to describe the heat transfer and desorption dynamics during secondary drying.(B) Structure of the proposed state observer. The observer receives the measured temperature and the control input (i.e., the shelf temperature and microwave power) at time step $t$, and estimates the temperature and concentration at the next time step. To initialize the observer at $t=0$, the estimated temperature can be set to the measured value. The initial concentration is completely unknown due to no measurement, so it can be set to some realistic values obtained from the literature or previous experiment.
@@
The most important consideration for observer design is to ensure that the estimated states converge to the true states fast compared to the time scale of the process, but have slow enough dynamics that the state estimates are insensitive to measurement noise. The time scales of the estimated states are specified by the observer gains $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$. Instead of attempting to search over all elements of the full observer gain matrices $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$, we show in Sections 3.1 and 3.2 that $\mathbf{L_{T}}$ and $\mathbf{L_{c}}$ be parameterized and rewritten as a product of the real scalar $L_{T}$ or $L_{c}$ and the matrix of ones, so the only design parameters are the real scalars $L_{T}$ and $L_{c}$.
The design procedure for the observer gains $L_{T}$ and $L_{c}$ consists of two main steps. The first step relies entirely on the mechanistic model and simulation, where the model prediction represents the true state of the system. In this step, a series of simulations are run for different values of the observer gains and parameters to investigate the dynamics of the system and observer under various conditions, and that information is used to select the observer gains based on the overall performance of the observer. The second step takes the real data/measurement from experiment into account, and so the observer gains are fine-tuned to be specific to the real system. These two steps are denoted as simulation-based observer design and experiment-based observer design, respectively. The former allows for different operating conditions and noise profiles to be tested so that the resulting design can cover many possible scenarios, whereas the latter is primarily specific to the system where the real data are available. This practical two-step procedure allows the observer to be designed efficiently without any complicated mathematical analysis, and is widely used in industrial applications. For the system, $L_{T}$ and $L_{c}$ have the units of s-1 and kg water/(kg solid $\cdot$ K $\cdot$ s), but only the magnitudes of $L_{T}$ and $L_{c}$ are reported to keep the plots simple and easy to visualize.
An alternative to the first step in observer design is to set the observer gains based on an analytical expression for the convergence time for the observer dynamics derived from the mathematical structure of the observer. To begin this analysis, first define the Jacobian of the nonlinear function $\mathbf{F}$ in (18) as
@@
where $\mathbf{F^{\prime}}$ can be calculated analytically or numerically. One of the simplest but efficient ways of analyzing a nonlinear state observer is to approximate the nonlinearities by linear equations, i.e., apply linearization. Linearization of the model equation (18) results in